OpenStudy (anonymous):

the endpoints of a diameter of a circle are (-7,3) and (5,1). what is the center of the circle?

OpenStudy (anonymous):

the midpoint of the line segment do you know how to find it?

OpenStudy (anonymous):

take the average in each coordinate \[(\frac{-7+5}{2},\frac{3+1}{2})\]

OpenStudy (jdoe0001):

\(\bf \textit{middle point of 2 points }\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ -7}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 5}}\quad ,&{\color{blue}{ 1}}) \end{array}\qquad \left(\cfrac{{\color{red}{ x_2}} + {\color{red}{ x_1}}}{2}\quad ,\quad \cfrac{{\color{blue}{ y_2}} + {\color{blue}{ y_1}}}{2} \right)\)